OpenStudy (anonymous):

OK i know there is 3 liters of juice , I need someone who could explain and help workout this equation PLEASE Harry has a jar that contains 75% of apple juice and the rest water. The bottle has 1 liter of water. Part A: Write an equation in one variable that can be used to find the total number of liters of apple juice and water in the jar. Define the variable used in the equation. (5 points) Part B: How many liters of apple juice are present in the jar? Show your work. (5 points)

OpenStudy (anonymous):

Ok do you know what a variable is?

OpenStudy (anonymous):

OK a variable is a letter representing a number which is not fully specified or is unknown

OpenStudy (anonymous):

So in this case you could use a as L variable. L would represent liters thats for number 1

OpenStudy (anonymous):

Harry has a jar that contains 75% of apple juice and the rest water. The bottle has 1 liter of water. Part A: Write an equation in one variable that can be used to find the total number of liters of apple juice and water in the jar. Define the variable used in the equation. Well you know apple juice = 75% so water = 25% Keep in mind that percentage is the number over 100. So when you have 75% of apple juice, it means you have 75/100 of apple juice from the total mixture of apple juice and water. The same applies for 25% water. Part B: How many liters of apple juice are present in the jar? Have x(or anything you want) represent the total volume of mixture. So you have for apple juice: 75/100 * x water: 25/100 * x You know the water is 1 liter. So you have 25/100 * x = 1 This is how you set up the equations. Can you do it from here? https://answers.yahoo.com/question/index?qid=20131009211325AAFKWDX

OpenStudy (anonymous):

if 25% = 1 liter, what does 75% =?

OpenStudy (anonymous):

My bad afk but yes i know what a variable is .. okay so know what from there?

OpenStudy (anonymous):

now*

OpenStudy (anonymous):

@sexyperson that uses 2 variables i can only use 1.

OpenStudy (anonymous):

well 3/4 of the jar is juice then 1/4 of the jar is water and there is 1 liter of water then there has to be 4 liters so juice would have to be 3 liters 3 liters of juice is in the jar

OpenStudy (anonymous):

let the total volume be x litres theres 75% apple juice (75/100)*x 0.75x total volume=0.75x+1 x=0.75x+1

OpenStudy (anonymous):

Thank you !!!!!!1

OpenStudy (anonymous):

OK so in the begining of the equation is says that 75% of the jar is apple juice and the rest is water. And a percent can only go to 100%. So that means 25% is water which is 1 liter. So there is 75% of the liquid liquid left and divide 75 by 25 and get three and that is how you know how many liters is apple juice.

OpenStudy (anonymous):

Sorry it took me so long

OpenStudy (anonymous):

Wrong

OpenStudy (anonymous):

How would you solve that?

OpenStudy (anonymous):

Objection

OpenStudy (anonymous):

He is way off

OpenStudy (anonymous):

okay im listening

OpenStudy (anonymous):

BAZINGA153760 lol

OpenStudy (anonymous):

You better be

OpenStudy (anonymous):

Okay so

OpenStudy (anonymous):

I was just making it a lot easier

OpenStudy (anonymous):

Spare me

OpenStudy (anonymous):

Let me solve this

OpenStudy (anonymous):

Lets see if you even know what you are doing

OpenStudy (anonymous):

Almost done, im just going to give her the answer. Because thats what she is looking for!

OpenStudy (anonymous):

lmfaoooooo

OpenStudy (anonymous):

Lets take a look at the first one

OpenStudy (anonymous):

Sorry im trying to help her understand how to do it her self instead of having to depend on people to give her the answer.

OpenStudy (anonymous):

Girl, dont make me snap on you!

OpenStudy (anonymous):

Okay so you explain, I give answer?

OpenStudy (anonymous):

ok and im not a girl

OpenStudy (anonymous):

Oh okay

OpenStudy (anonymous):

Explain and I will give da answer

OpenStudy (anonymous):

One person and I already gave an explanation OK

OpenStudy (anonymous):

OK SIR

OpenStudy (anonymous):

Let me take a gander

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

I dont sell myself for MEDALS

OpenStudy (anonymous):

I feel dirty

OpenStudy (anonymous):

Please take it back

OpenStudy (anonymous):

Cause you are . >.<

OpenStudy (anonymous):

What is that supposed to mean

OpenStudy (anonymous):

No fan either?

OpenStudy (anonymous):

Getting back on topic

OpenStudy (anonymous):

its the last one

OpenStudy (anonymous):

I already fanned you @Briaaa420

OpenStudy (anonymous):

Ill fan all of you

OpenStudy (anonymous):

I haven't even been paying attention

OpenStudy (anonymous):

You can take this one

OpenStudy (anonymous):

ok we shall be friends after this war of the brains @BAZINGA153760

OpenStudy (anonymous):

The sad thing is I dont even know what the problem is

OpenStudy (anonymous):

Friends , friends , i just want the answer outta all of this so please .

OpenStudy (anonymous):

Now she wants the answer

OpenStudy (anonymous):

WTF is going on here

OpenStudy (anonymous):

Idk you can take this one

OpenStudy (anonymous):

dakota , continue lol

OpenStudy (anonymous):

How do I get out of math

OpenStudy (anonymous):

OK so in the begining of the equation is says that 75% of the jar is apple juice and the rest is water. And a percent can only go to 100%. So that means 25% is water which is 1 liter. So there is 75% of the liquid liquid left and divide 75 by 25 and get three and that is how you know how many liters is apple juice.

OpenStudy (anonymous):

@BAZINGA153760 i have no idea

OpenStudy (anonymous):

i Know , i mentioned earlier i know there is 3 liters of juice , next?

OpenStudy (anonymous):

I'm so lost let me get one of my moderator fans to come help you k

OpenStudy (anonymous):

okay..

OpenStudy (anonymous):

@AravindG

OpenStudy (anonymous):

is coming to help you ok

OpenStudy (aravindg):

So is Part A done or should I start again?

OpenStudy (anonymous):

Start again pls

OpenStudy (anonymous):

ok thanks dakota!

OpenStudy (anonymous):

your welcome i'm not good at explaining things sorry i couldn't help:(

OpenStudy (aravindg):

From what I understood of the question we have a mixture of apple juice an water. Can you tell the total volume?

OpenStudy (anonymous):

4liters

OpenStudy (aravindg):

Find 75% of it you get amount of apple juice. (75/100)*4=x x is amount of apple juice. 4-x=Amount of water.

OpenStudy (anonymous):

Okay so how do i combine into only 1 equation...

OpenStudy (aravindg):

Sorry I think we made a mistake. The one liter given is in fact the 25% of mixture.So total volume is not 4 but a variable x. (25/100)x=1 By solving for total volume you get to calculate 75% of it which is amount of apple juice.

OpenStudy (anonymous):

x=0.75x+1 would this be the same? or noo

OpenStudy (aravindg):

0.75x+0.25x=x

OpenStudy (aravindg):

You need to solve for x from (25/100)x=1

OpenStudy (anonymous):

Okay now what would be the first step to solve?

OpenStudy (aravindg):

Then you can calculate 75% of it.